OpenStudy (anonymous):

Check my answer? Which best describes a system of equations that has no solution? 1. Consistent, independent 2. Inconsistent, dependent 3. inconsistant *

jimthompson5910 (jim_thompson5910):

you picked choice 3. inconsistent right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

Care to check a few more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

How many solutions does this system have? 2x+y=3 6x=9-3y I think it is infinite but you have 1, none, infinite or 2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

why do you think it's infinite?

jimthompson5910 (jim_thompson5910):

you are correct, hopefully you see how you're correct

OpenStudy (anonymous):

I do. I figured that was it because the other 3 were illogical. Thanks for the help!

jimthompson5910 (jim_thompson5910):

another way is to solve each for y 2x+y=3 y=3-2x y = -2x + 3

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

6x=9-3y 6x-9=-3y -3y = 6x - 9 y = (6x - 9)/(-3) y = (6x)/(-3) - 9/(-3) y = -2x + 3 so they are in fact the same line

OpenStudy (texaschic101):

jim is correct.....same line, infinite solutions